OpenStudy (anonymous):

Find the common solution. 8x + 13y = 2 y = x - 2 ( ? , ? )

OpenStudy (aroub):

lol how do u find the common solution??

OpenStudy (anonymous):

psh i dont know, thats why i asked(:

OpenStudy (aroub):

im so sorry i couldnt help u in this one!! :(

OpenStudy (anonymous):

its all good(: thank you though!

jhonyy9 (jhonyy9):

8x +13y =2 --- 13(x-2)=2-8x --- 13x -26=2-8x --- 21x = 28 --- x=28/21 = 4/3 y=x-2 --- y=4/3 -2=(4-6)/3 = -2/3 x=4/3 y= -2/3

OpenStudy (aroub):

ah , thank you jhonny9!! :D can u explain more please!, thank you! :)